What is the empirical formula for a compound which contains 0.0134 g of iron, 0.00769 g of sulfur and 0.0115 g or oxygen?
Answer #1

Fe mass = 0.0134 g

Fe molar mass = 55.84g

Fe moles = mass / molar mass = 0.0134 / 55.84 = 2.4 x 10^-4

S moles = 0.00769 / 32 = 2.4 x 10^-4

O moles = 0.0115 / 16 = 7.2 x 10^-4

mole ration of Fe , S ,and O

Fe : S : O = 2.4 x 10^-4   : 2.4 x 10^-4 : 7.2 x 10^-4

                 = 1 : 1: 3

so

Empirical formula = FeSO3